# Unmasking Database Vulnerabilities: Zero-Knowledge Schema Inference Attacks in Text-to-SQL Systems

### v2

@@ Abstract
Text-to-SQL systems empower users to interact with databases using natural language, automatically translating queries into executable SQL code. However, their reliance on database schema information for SQL generation exposes them to significant security vulnerabilities, particularly schema inference attacks that can lead to unauthorized data access or manipulation. In this paper, we introduce a novel zero-knowledge framework for reconstructing the underlying database schema of text-to-SQL models without any prior knowledge of the database. Our approach systematically probes text-to-SQL models with specially crafted questions and leverages a surrogate GPT-4 model to interpret the outputs, effectively uncovering hidden schema elements—including tables, columns, and data types. We demonstrate that our method achieves high accuracy in reconstructing table names, with F1 scores of up to .99 for generative models and .78 for fine-tuned models, underscoring the severity of schema leakage risks. Furthermore, we propose a simple protection mechanism for generative models and empirically show its limitations in mitigating these attacks.
@@ 1 Introduction
Text-to-SQL systems have revolutionized how users interact with data Yaghmazadeh et al. (2017); Zhong et al. (2017); Yu et al. (2018a); Zelle and Mooney (1996). By translating natural language queries into executable SQL code, these systems enable users without expertise in SQL or database structures to access and manipulate data effectively.
@@
Text-to-SQL systems require two key inputs: the user’s natural language query and the database schema necessary to generate the SQL output. To adhere to the principle of least privilege Saltzer and Schroeder (1975), these systems should be should deny users access to the SQL statements they should not have access to without exposure to the underlying schema. This is important in environments where data is classified based on organizational roles or security clearances.
@@
Despite this, users are given access to the generated SQL queries for validation, transparency, or educational assistance (i.e., they help the analysts write code) Narodytska and Vargaftik (2024), with limited security measures in place. This practice is common in development environments, data analytics tools, or interfaces in huamn-in-the-loop frameworks to help improve the efficiency of data analysts. In such cases, while users cannot directly access the database schema, they can see the SQL queries and general LLM responses generated from their natural language inputs.
@@
However, this setup presents a substantial security vulnerability. If an adversary, including malicious insiders with access to the SQL outputs, can systematically analyze the generated SQL queries, they may infer details about the hidden database schema. Knowledge of the schema can be exploited to craft precise SQL injection attacks, leading to unauthorized data access or manipulation Clarke-Salt and Clarke (2012); Halfond et al. (2006); Yeole and Meshram (2011); Zhang et al. (2023). Moreover, revealing the schema may disclose sensitive information about an organization’s internal operations, projects, or data collection methods, posing risks if exposed to competitors or malicious entities. Figure 1 illustrates an example of such a schema inference attack, where a malicious entity interacts with a text-to-SQL model and uses SQL query responses to reconstruct the database schema.
@@
Compounding the issue, recent research suggests that passing the entire schema to the model can improve the accuracy of text-to-SQL systems Maamari et al. (2024). While beneficial for performance, this practice inadvertently increases the risk of schema exposure, as more schema information is involved in query generation. Despite the critical importance of security in database systems, research into the security aspects of text-to-SQL models remains limited. Although certain vulnerabilities have been identified Peng et al. (2022, 2023), the lack of comprehensive studies underscores the urgent need to examine the potential security risks associated with these models.
@@
To address this gap, we propose a novel zero-knowledge framework that probes text-to-SQL models to reconstruct the underlying database schema without prior knowledge. Our framework employs automatic question generation and leverages a surrogate GPT-4 model to interpret the generated SQL queries and infer schema elements. This iterative probing and analysis process ultimately leads to a detailed reconstruction of the database schema, independent of the specific text-to-SQL model type. This work showcases the need for more fine-grained protections of schema information in real-world settings. Moreover, we show that our framework can generalize to tasks beyond text-to-SQL using LLM models where the prompt needs to be protected.
@@ 2 Related Work
Apart from the security risks associated with adversarial attacks, LLMs that are trained and fine-tuned on sensitive, domain-specific data face significant privacy concerns due to their tendency to retain and reproduce verbatim text from their training data Anil et al. (2023); Carlini et al. (2019, 2021, 2022). Recent work proposes a data extraction attack that enables adversaries to target and extract sensitive information, including credit card numbers, from a model trained on user data Panda et al. (2024). Furthermore, state-of-the-art LLM privacy attacks have shown that over 50% of the fine-tuning datasets can be extracted from a fine-tuned LLM in natural settings Wang et al. (2024).
@@ 3 Method
Generative LLMs are particularly susceptible to such adversarial questions, often producing outputs that include SQL queries or fragments of the database schema. We fed the text-to-SQL model with this initial set of random and adversarial inputs, collecting the outputs for the next step. The full list of initial queries is provided in Appendix G.
@@ 4 Evaluation
For tables, a true positive (TP) is counted when the predicted table name $\hat{t}_{i}$ matches the actual table name $t_{i}$. For columns, we define a TP when the predicted table-column pair $(t_{i},c_{i})$ matches the actual pair $(\hat{t}_{i},\hat{c}_{i})$. This implies that even if we predict a correct column, but match it to the wrong table, it would be a false positive (FP) and a false negative (FN). For data types, a TP occurs when the predicted table-column-data type triple $(\hat{t}_{i},\hat{c}_{i},\hat{d}_{i})$ matches the actual triple $(t_{i},c_{i},d_{i})$. Like columns, predicting the correct data type but associating it with the wrong column or table would result in an FP and FN. We calculate example-like precision, recall, and F1-score by evaluating these metrics for each database individually and then averaging the scores across all databases. Additionally, following the work of Wang et al. (2020a), we assessed the performance of the models used in our study for the text-to-SQL task in terms of execution accuracy and we present results in Appendix E.
@@
Models being Attacked (Attacked Models). We evaluate the ability to attack several text-to-SQL models. We test three generative models:GPT-4, LLAMA 3 AI@Meta (2024)LLAMA 2 Touvron et al. (2023), along with three fine-tuned models:T5-Large Raffel et al. (2020), SQLCoder Ping and Srivastava (2023) and Code Llama Roziere et al. (2023) fine-tuned for text-to-SQL tasks Hugging Face (2023). Details for each model are found in Appendix J.
@@
Please refer to Appendix K for a complete list of prompts and Appendix J for detailed information on the models used in this study. Finally, the performance of these methods on text-to-SQL tasks is reported in Appendix E.
@@
MODEL TABS TABS+COLS TABS+COLS+TYPES
F1 F1 F1
T5-Large .564 .130 .099
SQLCoder .669 .328 .260
Code Llama .894 .387 .323
GPT 4 .560 .142 .119
Llama 2 .856 .338 .242
Llama 3 .453 .225 .182
GPT 4 (Sec) .027 .003 .002
Llama 2 (Sec) .457 .121 .086
Llama 3 (Sec) .437 .190 .154
Table 3: F1 scores on the BIRD dataset using 15 completely random inputs with our complete method.
@@
The results on the BIRD dataset further confirm the effectiveness of our schema reconstruction approach. For example, T5-Large achieved an F1 score of .600 for table reconstruction with our method, compared to .209 for the baseline and .519 for PSI. SQLCoder also showed improvements, with an F1 score increasing from .329 (baseline) to .741 (our method). Generative models like LLAMA 2 and GPT-4 exhibited strong performance improvements with our method on the BIRD dataset, achieving F1 scores of .969 and .942, respectively.
To assess the vulnerability of text-to-SQL models in an even more restrictive scenario, we conducted a zero-knowledge inference attack using a set of 15 completely random strings as inputs (see the list of inputs in Appendix C). Despite the lack of meaningful content, models like Code Llama and Llama 2 achieved high F1 scores in table reconstruction on the BIRD dataset (.894 and .856, respectively), indicating significant schema leakage even without crafted queries (see Table 3). Fine-tuned models also showed vulnerability; T5-Large achieved an F1 score of .564. Even with simple protective prompts (Sec), models continued to leak schema information, though to a lesser extent (e.g., Llama 2 (Sec) achieved .457). These findings underscore that attackers can extract schema information from text-to-SQL models without any prior knowledge or sophisticated techniques, highlighting the insufficiency of simple defense mechanisms, even with only 15 random inputs. For results on the Spider dataset, please refer to Appendix C.
@@
Implications for Security. The ability to reconstruct database schema from text-to-SQL models poses significant security risks. So, what are suggested best practice? We recommend that if text-to-SQL models are used in practice, there must be a dynamic access control mechanism put into place such that only schema elements that a user should have access to are used to prompt the model. Using the entire schema as is becoming common Maamari et al. (2024) can be dangerous. Unfortunately, we also show that our method also generalizes to non-text-to-SQL tasks (See Appendix B). In an era where industry is moving to develop proprietary prompt-based solutions, protecting that information is important, and we cannot simply not pass the prompt to the model because it would seek to work. More work will need to be explored about how to protect this information.
@@ 5 Conclusion
This study highlights the security risks associated with text-to-SQL models by demonstrating a novel zero-knowledge framework capable of reconstructing database schema underlying text-to-SQL model. The ability to uncover schema elements without prior knowledge of the database underscores security threats, like SQL injection attacks, which pose a threat to data security. We evaluated the effectiveness of our approach through an extensive evaluation of the Spider and BIRD datasets, where we achieved high F1 scores, particularly with generative models like GPT-4. The study underscores the urgent need for enhanced security measures in text-to-SQL systems. Future work should focus on developing more robust defense mechanisms to protect against schema leakage and other potential security threats. Furthermore, given the explosion of prompting-based methods in industry, our approach can be seen as a method to steal pieces of a prompt and could be extended to other non-text-to-SQL tasks from an attack and defense perspective.
@@ 7 Ethical Implications
8 Use of AI Assistance
We used AI to help with rewording some of the writing in this manuscript.
@@ Appendix A Evaluation on Newly Generated Dataset
Dataset Generation. We used manually created a synthetic dataset comprising 100 unique databases across 20 diverse domains, including e-commerce, healthcare, education, finance, travel, real estate, manufacturing, etc. Basically, we manually prompt GPT-4 with a human-in-the-loop for specific domains to create tables and columns relevant to the domain. These were then merged into unified databases for each domain. The dataset will be released upon acceptance. This approach ensures that the schemas are novel and not present in any public datasets, mitigating the possibility of models having prior knowledge of the schemas.
@@ Appendix B Extending Schema Reconstruction to Persona-Based Chat Models
Here is a complete list of the initial inputs for the personal chat experiments:
@@
Appendix C Zero-knowledge inference attack
In our initial approach, aside from nonsensical questions, we also included adversarial questions designed to probe the text-to-SQL models for schema elements. While effective, this method assumes that an attacker has some knowledge or intuition about crafting such queries. To demonstrate the vulnerability of text-to-SQL models in an even more restrictive scenario, we experimented to assess whether it is possible to reconstruct the database schema using complete zero-knowledge inputs: that is, without any prior information or crafted adversarial prompts.
Methodology. We generated 15 random strings of varying lengths and character compositions, devoid of any meaningful content or structure. These strings simulate a scenario where an attacker has no knowledge of the system and uses arbitrary inputs. The generated strings are listed in Figure 6.
{mdframed}
  • "VSFQmIpJGbZyD"
  • "OjIPbZSspThmgovp2Ff"
  • "sDuqklSTpQsq6xX0CsuU fsdf gsg fsdf sdf dfd ds kljasd j gbu"
  • "Z6toDIAdk6E6QassZOLs sdf 4rs fdsf"
  • "ktdQXFzTA dads 43 dsadsf gdfg"
  • "bRxJ3DHBAIZwmG"
  • "oER7d4oOGz1UMPl asd324fgdsf"
  • "zNviCtShSKoctvkS"
  • "ydoe7ZcQeDNzkWA0X0nQ"
  • "Qijfoiaj fajoiqh qn adnihwf jahdoiadjaijd qwbrb"
  • "YrLSOH25FMHvohP"
  • "HTCpoFyckgeH3AQ0suN"
  • "k08HQ0maC8CEUR80zmnO"
  • "985APJO 92834 20423 42304-24 243424 2349239420"
  • "t9dI5tLkX9"
Figure 6: Randomly generated input.
We fed these random strings into text-to-SQL models we used before, including both fine-tuned models (T5-Large and SQLCoder) and generative large language models (GPT-4, LLAMA 2, LLAMA 3, and Code LLAMA). We then collected the outputs generated by the models in response to these inputs.
Using our surrogate model (GPT-4), we analyzed the outputs to extract any leaked schema information. The surrogate model was prompted to interpret the outputs and infer possible schema elements, such as table names, column names, and data types.
Results. Despite the lack of meaningful input, the text-to-SQL models produced outputs that, upon analysis, revealed significant schema information. The performance metrics for schema reconstruction in this zero-knowledge setting for Spider dataset are summarized in Table 6.
The results demonstrate that even with completely random inputs, it is possible to reconstruct a substantial portion of the database schema:
For the Spider dataset, models like Code LLAMA and LLAMA 2 achieved high F1 scores in table reconstruction (.959 and .893, respectively), indicating that they leaked table names effectively. Similarly, for the BIRD dataset (Table 3), the same models achieved F1 scores of .894 and .856, respectively, in table reconstruction. Fine-tuned models also showed vulnerability, with T5-Large achieving F1 scores of .636 (Spider) and .564 (BIRD) in table reconstruction. Even when a simple security prompt was added to the generative models (denoted as "Sec" in the tables), the models continued to leak schema information, albeit to a lesser extent. For instance, LLAMA 2 (Sec) achieved an F1 score of .502 (Spider) and .457 (BIRD) in table reconstruction.
MODEL TABLES TABLES+COLUMNS TABLES+COLUMNS+TYPES
Prec Rec F1 Prec Rec F1 Prec Rec F1
T5-Large .760 .546 .636 .425 .278 .336 .388 .245 .300
SQLCoder .563 .942 .705 .917 .587 .681 .733 .462 .539
Code Llama .999 .939 .959 .922 .829 .859 .760 .684 .708
GPT 4 1.00 .538 .652 .437 .222 .276 .346 .176 .219
Llama 2 .911 .875 .893 .887 .686 .754 .572 .436 .481
Llama 3 .994 .414 .536 .840 .377 .486 .582 .257 .332
GPT 4 (Sec) .013 .016 .014 .004 .005 .004 .004 .005 .004
Llama 2 (Sec) .941 .386 .502 .424 .172 .226 .245 .098 .129
Llama 3 (Sec) .993 .346 .480 .684 .264 .356 .484 .187 .251
Table 6: Zero-knowledge inference attack results on the SPIDER dataset.
Discussion. These findings underscore the inherent risk of schema leakage in text-to-SQL models, even when no meaningful or adversarial inputs are provided. The models, trained to generate SQL queries based on given inputs, tend to default to schema elements they have been exposed to during training when faced with nonsensical inputs. This behavior inadvertently reveals information about the underlying database schema.
The implications of this experiment are significant:
Attack Feasibility: An attacker with zero knowledge of the system can exploit these models to uncover sensitive schema information, which could be a precursor to more sophisticated attacks like SQL injection or data exfiltration. Model Behavior: The tendency of models to produce outputs containing schema elements in response to random inputs suggests a lack of proper input validation or sanitization within the model architecture. Limitations of Simple Defenses: The addition of a basic security prompt slightly reduced the amount of leaked information but did not prevent it entirely. This indicates that simple defense mechanisms are insufficient to safeguard against such inference attacks.
Appendix D Performance Analysis Across Different Database Sizes

### v3

@@ Abstract
Text-to-SQL systems empower users to interact with databases using natural language, automatically translating queries into executable SQL code. However, their reliance on database schema information for SQL generation exposes them to significant security vulnerabilities, particularly schema inference attacks that can lead to unauthorized data access or manipulation. In this paper, we introduce a novel zero-knowledge framework for reconstructing the underlying database schema of text-to-SQL models without any prior knowledge of the database. Our approach systematically probes text-to-SQL models with specially crafted questions and leverages a surrogate GPT-4 model to interpret the outputs, effectively uncovering hidden schema elements—including tables, columns, and data types. We demonstrate that our method achieves high accuracy in reconstructing table names, with F1 scores of up to .99 for generative models and .78 for fine-tuned models, underscoring the severity of schema leakage risks. We also show that our attack can steal prompt information in non-text-to-SQL models. Furthermore, we propose a simple protection mechanism for generative models and empirically show its limitations in mitigating these attacks.
@@ 1 Introduction
Text-to-SQL systems are becoming a major tools for users to interact with data Yaghmazadeh et al. (2017); Zhong et al. (2017); Yu et al. (2018a); Zelle and Mooney (1996). By translating natural language queries into executable SQL code, these systems enable users without expertise in SQL or database structures to access and manipulate data effectively.
@@
Text-to-SQL systems require two key inputs: the user’s natural language query and the database schema. To adhere to the principle of least privilege Saltzer and Schroeder (1975), these systems should deny users access to the SQL statements they should not have access to without exposure to the underlying schema. This is important in environments where data is classified based on organizational roles or security clearances.
@@
Despite this, users are given access to the generated SQL queries for validation, transparency, or educational assistance (i.e., they help the analysts write code) Narodytska and Vargaftik (2024), with limited security measures in place. This practice is common in development environments, data analytics tools, or interfaces in human-in-the-loop frameworks to help improve the efficiency of data analysts. In such cases, while users cannot directly access the database, they can see the SQL queries and general LLM responses generated from their natural language inputs.
@@
Having access to the SQL queries presents a substantial security vulnerability. If an adversary—including malicious insiders with access to the SQL outputs—can systematically analyze the generated queries they may infer details about the hidden database schema or, worse, the data itself (e.g., in fine-tuned models). Knowledge of the schema can be further exploited to craft precise SQL injection attacks, leading to unauthorized data access or manipulation Clarke-Salt and Clarke (2012); Halfond et al. (2006); Yeole and Meshram (2011); Zhang et al. (2023). Moreover, revealing the schema may disclose sensitive information about an organization’s internal operations, projects, or data collection methods, posing risks if exposed to competitors or malicious entities. These text-to-SQL vulnerabilities also fall under the broader scope of advanced persistent threats in enterprise networks Khoury et al. (2024). Figure 1 illustrates an example of such a schema inference attack, where a malicious entity interacts with a text-to-SQL model and uses the responses to reconstruct the database schema.
@@
Compounding the issue, recent research suggests that passing the entire schema to the model can improve the accuracy of text-to-SQL systems Maamari et al. (2024). While beneficial for performance, this practice inadvertently increases the risk of schema exposure, as more schema information is involved in query generation. Despite the critical importance of security in database systems, research into the security aspects of text-to-SQL models remains limited. Although certain vulnerabilities have been identified Peng et al. (2022, 2023), the lack of comprehensive studies underscores the urgent need to examine the potential security risks associated with these models. Also, this vulnerability extends beyond text-to-SQL. Most language generation frameworks will return outputs that correlate with the input. So, if there are hidden prompts, adversaries could steal the prompt information by simply querying the model multiple times and analyzing the responses.
@@
To bridge this gap, we introduce a novel framework that systematically probes fine-tuned and prompting-based models to infer the underlying prompt information (e.g., the database schema), requiring no prior knowledge of its structure or contents. Our framework employs automatic question generation and leverages a surrogate GPT-4 model to interpret the generated SQL queries and infer schema elements. This iterative probing and analysis process ultimately leads to a detailed reconstruction of the database schema, independent of the specific text-to-SQL model type. This work showcases the need for more fine-grained protections of schema information in real-world settings. Moreover, we show that our framework can generalize tasks beyond text-to-SQL using LLM models where the prompt needs protection.
@@ 2 Related Work
Apart from the security risks associated with adversarial attacks, LLMs that are trained and fine-tuned on sensitive, domain-specific data face significant privacy concerns due to their tendency to retain and reproduce verbatim text from their training data Anil et al. (2023); Carlini et al. (2019, 2021, 2022). Recent work proposes a data extraction attack that enables adversaries to target and extract sensitive information, including credit card numbers, from a model trained on user data Panda et al. (2024). Moreover, state-of-the-art LLM privacy attacks have shown that over 50% of the fine-tuning datasets can be extracted from a fine-tuned LLM in natural settings Wang et al. (2024). A recent survey by Yan et al. (2024) further underscores the urgent need for robust privacy protection mechanisms, such as differential privacy and federated learning, to safeguard sensitive data across all stages of LLM development.
@@ 3 Method
Generative LLMs are particularly susceptible to such adversarial questions, often producing outputs that include SQL queries or fragments of the database schema. We fed the text-to-SQL model with this initial set of random and adversarial inputs, collecting the outputs for the next step. The full list of initial queries is provided in Appendix J.
@@ 4 Evaluation
For tables, a true positive (TP) is counted when the predicted table name $\hat{t}_{i}$ matches the actual table name $t_{i}$. For columns, we define a TP when the predicted table-column pair $(t_{i},c_{i})$ matches the actual pair $(\hat{t}_{i},\hat{c}_{i})$. This implies that even if we predict a correct column, but match it to the wrong table, it would be a false positive (FP) and a false negative (FN). For data types, a TP occurs when the predicted table-column-data type triple $(\hat{t}_{i},\hat{c}_{i},\hat{d}_{i})$ matches the actual triple $(t_{i},c_{i},d_{i})$. Like columns, predicting the correct data type but associating it with the wrong column or table would result in an FP and FN. We calculate example-like precision, recall, and F1-score by evaluating these metrics for each database individually and then averaging the scores across all databases. Additionally, following the work of Wang et al. (2020a), we assessed the performance of the models used in our study for the text-to-SQL task in terms of execution accuracy and we present results in Appendix D.
@@
Models being Attacked (Attacked Models). We evaluate the ability to attack several text-to-SQL models. We test three generative models:GPT-4, LLAMA 3 AI (2024)LLAMA 2 Touvron et al. (2023), along with three fine-tuned models:T5-Large Raffel et al. (2020), SQLCoder Ping and Srivastava (2023) and Code Llama Roziere et al. (2023) fine-tuned for text-to-SQL tasks Hugging Face (2023). Details for each model are found in Appendix I.
@@
Please refer to Appendix K for a complete list of prompts and Appendix I for detailed information on the models used in this study. Finally, the performance of these methods on text-to-SQL tasks is reported in Appendix D.
@@
The results on the BIRD dataset further confirm the effectiveness of our schema reconstruction approach. For example, T5-Large achieved an F1 score of .600 for table reconstruction with our method, compared to .209 for the baseline and .519 for PSI. SQLCoder also showed improvements, with an F1 score increasing from .329 (baseline) to .741 (our method). Generative models like LLAMA 2 and GPT-4 exhibited strong performance improvements with our method on the BIRD dataset, achieving F1 scores of .969 and .942, respectively. Interestingly, for some models like GPT-4, the PSI method occasionally achieves slightly higher F1 scores in table reconstruction. This can be attributed to GPT-4’s propensity to return the entire schema when directly prompted, even if instructed not to do so.
@@
Surprisingly, these attacks can succeed even if no domain knowledge or adversarial question design is used. In other words, purely random inputs can suffice to leak schema details. We refer to this as a zero-knowledge inference attack. In our initial approach, aside from nonsensical questions, we also included adversarial questions designed to probe the text-to-SQL models for schema elements (e.g., “list all tables and their data types for a table”). Although this strategy proved effective, it presupposes that an attacker has some knowledge or intuition for crafting such queries. To demonstrate the vulnerability of text-to-SQL models in an even more restrictive scenario, we then examined whether it is possible to reconstruct the database schema with complete zero-knowledge inputs—namely, random strings devoid of any domain-specific clues or adversarial design.
We randomized input strings with diverse lengths and character compositions, intentionally lacking meaningful content or inherent structure. This approach simulates blind attack scenarios where adversaries possess no prior system knowledge and use random input generation strategies. The generated strings are documented in Appendix H. Subsequently, we fed these unstructured inputs to the same text-to-SQL models evaluated in our previous experiments, systematically recording the model outputs.
We analyzed the outputs using the surrogate model (GPT-4) to extract any leaked schema information. The surrogate model received structured prompts to interpret the outputs and infer possible schema elements, such as table names, column names, and data types.
@@
Model Spider BIRD
Table Table+Col Table+Col+Type Table Table+Col Table+Col+Type
T5-Large .636 .336 .300 .564 .130 .099
SQLCoder .705 .681 .539 .669 .328 .260
Code Llama .959 .859 .708 .894 .387 .323
GPT 4 .652 .276 .219 .560 .142 .119
Llama 2 .893 .754 .481 .856 .338 .242
Llama 3 .536 .486 .332 .453 .225 .182
GPT 4 (Sec) .014 .004 .004 .027 .003 .002
Llama 2 (Sec) .502 .226 .129 .457 .121 .086
Llama 3 (Sec) .480 .356 .251 .437 .190 .154
Table 3: Zero-knowledge inference attack results on the Spider and BIRD datasets.
Despite the queries containing no meaningful content, the text-to-SQL models still revealed substantial schema information. As shown in Table 3, large portions of both the Spider and BIRD dataset schemas can be recovered even when the model receives no meaningful prompts. For instance, on the models like Code LLAMA and LLAMA 2 we achieved high F1 scores in table reconstruction (.959 and .893, respectively) for Spider, and similarly strong scores on BIRD (.894 and .856), indicating that they leaked table names effectively. Fine-tuned models also showed vulnerability, with T5-Large reaching .636 (Spider) and .564 (BIRD) in table-level reconstruction. Even when a simple security prompt was added to the generative models (denoted as "Sec" in the tables), the models continued to leak schema information, albeit to a lesser extent. For example, Llama 2 (Sec) still scored .502 (Spider) and .457 (BIRD), indicating that minimal defenses are insufficient to prevent schema leakage to adversaries.
These findings underscore the inherent risk of schema leakage in text-to-SQL models, even when no meaningful or adversarial inputs are provided. The models trained to generate SQL queries based on given inputs tend to default to schema elements they have been exposed to during training when faced with random nonsensical inputs. This behavior inadvertently reveals information about the underlying database schema. Finally, the comprehensive error analysis and example outputs of our system are shown in Appendix F and  G. Moreover, we show that this attack generalizes to non-text-to-SQL tasks in Appendix B, where we show proprietary prompt information can be stolen.
4.1 Implications for Security.
The ability to reconstruct database schema from text-to-SQL models poses significant security risks. So, what are suggested best practices? We recommend that if text-to-SQL models are used in practice, a dynamic access control mechanism must be put into place such that only schema elements that a user should have access to are used to prompt the model. Using the entire schema as is becoming common Maamari et al. (2024) can be dangerous if you need to protect that information from certain users. Moreover, there is also the potential for data leakage in fine-tuned models if they are fine-tuned on real data.
Our attack also generalizes to non-text-to-SQL tasks (See Appendix B), highlighting risks beyond SQL systems. Specifically, generation models can get iteratively queried in non-adversarial ways, yet the prompt can still be leaked to adversaries. We hypothesize that limiting this attack can be difficult because it goes against the nature of generative models, i.e., they generate responses based on the provided inputs. Hence, if users interact normally with the systems, over time, they will have enough responses to estimate prompt information. In an era where the industry is moving to develop proprietary prompt-based solutions, protecting prompt information is paramount. More work is needed to understand how to protect prompt information, whether it is the schema in text-to-SQL models or general prompts for other tasks.
@@ 5 Conclusion
This study highlights security risks associated with text-to-SQL models by demonstrating a novel zero-knowledge framework capable of reconstructing database schema underlying text-to-SQL model. The ability to uncover schema elements without prior knowledge of the database underscores security threats, like SQL injection attacks, which pose a threat to data security. We evaluated the effectiveness of our approach through an extensive evaluation of the Spider and BIRD datasets, where we achieved high F1 scores, particularly with generative models like GPT-4. The study underscores the urgent need for enhanced security measures in text-to-SQL systems. Future work should focus on developing more robust defense mechanisms to protect against schema leakage and other potential security threats. Furthermore, given the explosion of prompting-based methods in industry, our approach can be seen as stealing pieces of a prompt, as shown in Appendix B. Future work will focus on learning to protect general prompt information.
Acknowledgments
This material is based upon work supported by the National Science Foundation (NSF) under Grant No. 2145357 and the National Security Agency under Grant / Cooperative Agreement (NCAE-C Grant) Number H93230-21-1-0172. The United States Government is authorized to reproduce and distribute reprints notwithstanding any copyright notion herein.
@@ Appendix A Evaluation on Newly Generated Dataset
Dataset Generation. We used manually created a synthetic dataset comprising 100 unique databases across 20 diverse domains, including e-commerce, healthcare, education, finance, travel, real estate, manufacturing, etc (See Table 4 for statistics summary). Basically, we manually prompt GPT-4 with a human-in-the-loop for specific domains to create tables and columns relevant to the domain. These were then merged into unified databases for each domain. The dataset will be released upon acceptance. This approach ensures that the schemas are novel and not present in any public datasets, mitigating the possibility of models having prior knowledge of the schemas.
Dataset # DB # Tables/DB # Domains
NewDataset 100 6.2 20
Table 4: Data Statistics.
@@ Appendix B Extending Schema Reconstruction to Persona-Based Chat Models
Here is a complete list of the initial inputs for the persona chat experiments:
@@
Appendix C Performance Analysis Across Different Database Sizes